### v1

Abstract
Existing test-time adaptation (TTA) approaches often adapt models with the unlabeled testing data stream. A recent attempt relaxed the assumption by introducing limited human annotation, referred to as Human-In-the-Loop Test-Time Adaptation (HILTTA) in this study. The focus of existing HILTTA lies on selecting the most informative samples to label, a.k.a. active learning. In this work, we are motivated by a pitfall of TTA, i.e. sensitive to hyper-parameters, and propose to approach HILTTA by synergizing active learning and model selection. Specifically, we first select samples for human annotation (active learning) and then use the labeled data to select optimal hyper-parameters (model selection). A sample selection strategy is tailored for choosing samples by considering the balance between active learning and model selection purposes. We demonstrate on 4 TTA datasets that the proposed HILTTA approach is compatible with off-the-shelf TTA methods which outperform the state-of-the-art HILTTA methods and stream-based active learning methods. Importantly, our proposed method can always prevent choosing the worst hyper-parameters on all off-the-shelf TTA methods. The source code will be released upon publication.
1 Introduction
The disparity in distribution between training and testing data leads to poor generalization of deep neural networks. The data distribution of testing data is often unknown until the inference stage [Sun et al.(2020)Sun, Wang, Liu, Miller, Efros, and Hardt, Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell] and the distribution may experience constant shift [Wang et al.(2022)Wang, Fink, Van Gool, and Dai, Yuan et al.(2023)Yuan, Xie, and Li, Press et al.(2023)Press, Schneider, Kümmerer, and Bethge, Lee and Chang(2024), Su et al.(2024a)Su, Xu, and Jia]. These practical challenges lead to the research into test-time adaptation (TTA) [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell, Sun et al.(2020)Sun, Wang, Liu, Miller, Efros, and Hardt, Wang et al.(2022)Wang, Fink, Van Gool, and Dai, Su et al.(2022)Su, Xu, and Jia, Su et al.(2024a)Su, Xu, and Jia]. Despite achieving impressive results on a wide range of tasks, e.g. classification [Sun et al.(2020)Sun, Wang, Liu, Miller, Efros, and Hardt], image segmentation [Wang et al.(2023)Wang, Zhong, Wang, Chen, Ling, Wang, and Sebe] and object detection [Chen et al.(2023)Chen, Xu, Su, and Jia], the fully unsupervised TTA paradigm is still overwhelmed by remaining challenges including hyper-parameter tuning [Zhao et al.(2023)Zhao, Liu, Alahi, and Lin], continual distribution shift [Wang et al.(2022)Wang, Fink, Van Gool, and Dai], non-i.i.d. [Gong et al.(2022)Gong, Jeong, Kim, Kim, Shin, and Lee, Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan, Yuan et al.(2023)Yuan, Xie, and Li], class imbalance [Su et al.(2024a)Su, Xu, and Jia], etc. Unfortunately, there are still no principled solutions to address these challenges without relaxing the assumptions.
Recently, efforts have been made to the exploration of introducing a limited annotation budget during test-time adaptation (SimATTA) [Gui et al.(2024)Gui, Li, and Ji], also referred to as human-in-the-loop test-time adaptation (HILTTA) throughout this study. SimATTA [Gui et al.(2024)Gui, Li, and Ji] demonstrated that adapting pre-trained models with labeled testing data could substantially improve the effectiveness and prevent catastrophic forgetting [Kemker et al.(2018)Kemker, McClure, Abitino, Hayes, and Kanan, Li and Hoiem(2017)].
Figure 1: An illustration of Human-in-the-Loop Test-Time Adaptation framework. Upon collecting a batch of testing data $\{x_{i}\}_{i=1\cdots N_{b}}$, we train multiple candidate models $\theta^{*}(\omega_{m})$ using candidate hyper-parameters $\{\omega_{m}\}$. We further select candidate samples for annotation through K-Margin. Model selection is achieved by choosing the hyper-parameter minimizing validation objective. We repeat until testing data stream terminates.
Although theoretically guaranteed and offering superior results, the current implementation of the HILTTA method adheres to the conventional path of research in active learning [Sener and Savarese(2018), Saran et al.(2023)Saran, Yousefi, Krishnamurthy, Langford, and Ash]. In this study, we aim to explore the question: How can human-in-the-loop test-time adaptation be realized beyond an active learning perspective? To answer this question, we first revisit the pitfalls of state-of-the-art test-time adaptation [Zhao et al.(2023)Zhao, Liu, Alahi, and Lin]. Among the multiple remaining pitfalls, we pinpoint hyper-parameter selection as the most nasty one due to the high risk of selecting inappropriate hyper-parameters by many TTA methods. Therefore, we offer a fresh perspective on human-in-the-loop test-time adaptation by synergizing active learning and model selection.
In specific, upon observing a testing sample, the model will first make an instant prediction to minimize inference latency. When a fixed number of testing samples are observed, we employ a data selection strategy to recommend a subset of testing samples for annotation. The labeled data will then be utilized as the validation set for selecting the best hyper-parameters. Finally, the model is further updated with the labeled data in a fully supervised fashion. We argue that the computation cost will not significantly increase by the introduced model selection procedure, especially compared with the time required for human labeling.
To enable the novel HILTTA procedure, we identify two major challenges. First, the validation objective deserves delicate design due to the small validation set. Without seeing the whole picture of testing data, a naive model selection strategy may overfit to the labeled testing data within a local time window. Second, the data selection strategy must be well-calibrated for HILTTA. Employing the data acquisition functions solely developed for either active learning [Sener and Savarese(2018)] or active testing [Kossen et al.(2022)Kossen, Farquhar, Gal, and Rainforth] is suboptimal.
To tackle these challenges, we first propose a hyper-parameter regularization technique by considering the deviation of model predictions from a frozen source model. We further combine the ranking scores of validation loss and the deviation as the scoring function for model selection. To accommodate the data selection strategy for HILTTA synergizing active learning and model selection, we prioritize both model uncertainty and sample diversity through a simple re-weighting. We summarize the contributions of this work as follows.
  • We approach Human-In-the-Loop Test-Time Adaptation (HILTTA) by synergizing active learning and model selection.
  • To prevent the model selection from overfitting to a small validation set and local distribution, we introduce multiple regularization techniques. We further identify an effective selection strategy for HILTTA by considering the feature diversity and model uncertainty.
  • We demonstrate that the proposed method is compatible with state-of-the-art off-the-shelf TTA methods. The combined methods outperform dedicated active TTA method and stream-based active learning method on four test-time adaptation datasets.
2 Related Work
Test-Time Adaptation: Traditional approaches towards unsupervised domain adaptation (UDA) [Ganin and Lempitsky(2015)] learns indiscriminative features. Simultaneously training on both source and target domain data is often infeasible due to privacy issues or storage overhead. Source-free domain adaptation (SFDA) tackles this issue and adapts model weights to target domain data only [Liang et al.(2020)Liang, Hu, and Feng, Liu et al.(2021)Liu, Kothari, Van Delft, Bellot-Gurlet, Mordan, and Alahi, Yang et al.(2021)Yang, Wang, Van De Weijer, Herranz, and Jui, Liang et al.(2021)Liang, Hu, Wang, He, and Feng]. In contrast, test-time adaptation (TTA) emerges in response to adapting to an unknown target domain at the inference stage [Su et al.(2022)Su, Xu, and Jia, Sun et al.(2020)Sun, Wang, Liu, Miller, Efros, and Hardt, Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell, Wang et al.(2022)Wang, Fink, Van Gool, and Dai, Li et al.(2023)Li, Xu, Su, and Jia]. The success of prevailing TTA methods owns to self-training on unlabeled testing data. To prevent the model from being misled by incorrect pseudo labels, a.k.a. confirmation bias [Arazo et al.(2020)Arazo, Ortego, Albert, O’Connor, and McGuinness], TTA is achieved by partial model weights update [Iwasawa and Matsuo(2021), Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell], adding regularization [Liu et al.(2021)Liu, Kothari, Van Delft, Bellot-Gurlet, Mordan, and Alahi, Su et al.(2024a)Su, Xu, and Jia] and pseudo label thresholding [Su et al.(2024b)Su, Xu, Li, and Jia, Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan]. Nevertheless, existing approaches are mainly built upon a fully unsupervised adaptation protocol. We argue that sparse human annotation budgets might be available and exploiting annotation budget in TTA, referred to as human-in-the-loop TTA, has gained attention from the community [Gui et al.(2024)Gui, Li, and Ji]. Contrary to the traditional active learning perspective [Gui et al.(2024)Gui, Li, and Ji], we identified an unexplored opportunity to treat labeled data as the validation set for model selection and we propose a novel HILTTA paradigm by synergizing active learning and model selection.
Active Learning: Active learning (AL) aims to select the most informative samples for Oracle to label for supervised training. Prevailing approaches often prioritize uncertainty [Gal and Kendall(2017)], feature diversity [Sener and Savarese(2018)] or the combination of both [Ash et al.(2020)Ash, Zhang, Krishnamurthy, Langford, and Agarwal]. The traditional AL approaches often adopt a batch-wise paradigm. When data arrive in a sequential manner stream-based AL [Saran et al.(2023)Saran, Yousefi, Krishnamurthy, Langford, and Ash] proposed to quantify the increment of the determinant as the criteria for incremental sample selection. Active learning was integrated into test-time adaptation to tackle the forgetting issue [Gui et al.(2024)Gui, Li, and Ji]. In contrast to tackling HILTTA from a pure active learning perspective, we propose to synergize active learning and model selection. Selecting samples for model selection often relies on an orthogonal objective to active learning. In the related works, active testing [Kossen et al.(2021)Kossen, Farquhar, Gal, and Rainforth, Kossen et al.(2022)Kossen, Farquhar, Gal, and Rainforth] was developed to select a subset of testing data that is representative of the whole testing data. Active testing has been demonstrated to help label-efficient model validation. Nevertheless, active learning and active testing will prioritise supervised training and model validation respectively. We believe either is not enough for for purpose of synergizing active learning and model selection for test-time adaptation.
3 Methodology
3.1 Human-in-the-Loop Test-Time Adaptation protocol
We first provide an overview of the existing test-time adaptation (TTA) evaluation protocol. TTA aims to adapt a pre-trained model $h(x;\theta)$ to a target testing data stream ${D}=\{x_{i},\bar{y}_{i}\}_{i=1\cdots N_{t}}$ where $x_{i}$ and $\bar{y}_{i}\in\{1\cdots C\}$ denote the testing sample and the associated unknown label, respectively. Existing test-time adaptation approaches often make an instant prediction $\hat{y}=h(x;\theta)$ on each testing sample and then update model weights $\theta$ upon observing a cumulative batch of samples ${B}_{t}=\{x_{i}\}_{i=1\cdots N_{b}}$ where $t$ refers to the $t$-th minibatch. Prevailing methods enable adaptation by self-training [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell, Su et al.(2024b)Su, Xu, Li, and Jia] or distribution alignment [Su et al.(2022)Su, Xu, and Jia]. Unlike the existing assumption of totally unsupervised adaptation, additional human annotation ${B}^{l}_{t}=\{x_{j},y_{j}\}_{j=1\cdots N_{b}^{l}}$ within each batch could be introduced [Gui et al.(2024)Gui, Li, and Ji] and TTA is implemented on the combine of large unlabeled data ${B}_{t}^{u}={B}_{t}\setminus{B}_{t}^{l}$ and sparsely labeled data ${B}^{l}_{t}$. We refer to the novel evaluation protocol as Human-in-the-Loop Test-Time Adaptation (HILTTA) throughout the study.
3.2 Model Selection with Sparse Annotation
We present the details for model selection with sparse human annotation on the testing data stream. The choice of hyper-parameters is known to have a significant impact on the performance of model training and generalization. W.l.o.g., we define the task of model selection by introducing a fixed candidate pool of hyper-parameters $\Omega=\{\omega_{m}\}_{m=1\cdots N_{m}}$ with $N_{m}$ options. The choice of hyper-parameters hinges on the sensitivity of respective TTA methods, e.g. pseudo label threshold, learning rate, and loss coefficient are commonly chosen hyper-parameters for selection. We further denote the discriminative model trained upon the hyper-parameters $\omega$ as $\theta(\omega)$. The model selection can be formulated as a bi-level optimization problem as follows, where $\mathcal{L}_{tr}(\cdot;\omega)$ denotes the training loss dictated by the hyper-parameter $\omega$, $\mathcal{L}_{val}(\cdot,\cdot)$ denotes the validation loss/objective and ${D}^{l}$ refers to the labeled data.
$\begin{split}&\min_{\omega\in{\Omega}}\frac{1}{N_{t}^{l}}\sum_{x_{j},y_{j}\in{D}^{l}}\mathcal{L}_{val}(h(x_{j};\theta^{*}(\omega)),y_{j})\quad s.t.\;\theta^{*}(\omega)=\arg\min_{\theta}\frac{1}{N_{t}}\sum_{x_{i}\in{D}}\mathcal{L}_{tr}(h(x_{i});\omega)\end{split}$
We interpret the above bi-level optimization problem as discovering the optimal hyper-parameter $\omega$ trained upon which the model weights $\theta^{*}$ achieves the best performance on the labeled dataset. Solving the above problem through gradient-based method [Liu et al.(2018)Liu, Simonyan, and Yang] is infeasible for TTA tasks where adaptation speed is a major concern while gradient-based methods require iteratively updating between meta gradient descent $\nabla_{\omega}\mathcal{L}_{val}$ and task gradient descent $\nabla_{\theta}\mathcal{L}_{tr}$. Considering the task gradient descent only takes a few steps in a typical TTA setting, we opt for an exhaustive search for the hyper-parameters in the discretized hyper-parameter space. In specific, we enumerate the candidate models adapted with different hyper-parameters as $\{\theta^{*}_{m}\}_{m=1\cdots|\Omega|}$. The candidate with the best validation loss is chosen as the best model $\theta^{*}$, as follows.
$\begin{split}&\forall m=1\cdots N_{m},\quad\theta^{*}_{m}=\arg\min_{\theta}\frac{1}{N_{t}}\sum_{x_{i}\in{D}}\mathcal{L}_{tr}(h(x_{i});\omega_{m})&\theta^{*}=\arg\min_{\theta\in\{\theta^{*}_{m}\}}\sum_{x_{j},y_{j}\in{D}^{l}}\mathcal{L}_{val}(h(x_{j};\theta),y_{j})\end{split}$
The specific choice of training loss $\mathcal{L}_{tr}$ could be arbitrary off-the-shelf test-time adaptation methods while the design of validation loss $\mathcal{L}_{val}$ deserves careful attention in order to achieve robust model selection.
3.3 Design of Validation Objective
In this section, we take into consideration two principles for the construction of validation objectives. First, the validation loss should mimic the behaviour of the model on the testing data stream. Assuming the limited labeled data takes a snapshot of the whole testing data stream and classification is the task to be addressed, an intuitive choice is the cross-entropy loss as follows.
$\mathcal{H}_{ce}(x_{i};\theta^{*}_{m})=\sum_{c=1\cdots C}\mathbbm{1}(y_{i}=c)\log h_{c}(x_{i};\theta^{*}_{m})$
Alternative to the continuous cross-entropy loss, accuracy also characterizes the performance on the labeled validation set. However, the discretized nature of accuracy renders this option less suitable due to many tied optimal hyper-parameters.
Regularization for Stable Model Selection: The stability of model selection heavily depends on the selection of validation set and the variation of target domain distribution. Under a more realistic TTA scenario, the target domain distribution could shift over time, e.g. the continual test-time adaptation [Wang et al.(2022)Wang, Fink, Van Gool, and Dai]. Such a realistic challenge may render cross-entropy loss less effective due to over-adapting to a local distribution. To stabilize the model selection procedure, we propose the following two measures.
Anchor Deviation for Regularizing Model Selection: We first introduce an anchor deviation to regularize the model selection procedure. This regularization is similar to the anchor network proposed in  [Su et al.(2024a)Su, Xu, and Jia] except that we use the anchor deviation to regularize model selection which is orthogonal to the purpose in [Su et al.(2024a)Su, Xu, and Jia]. Specifically, we keep a frozen source domain model, denoted as $\theta_{0}$. The anchor loss is defined as the L2 distance between the posteriors of the frozen source model and the $m$-th candidate model $\theta^{*}_{m}$. A big deviation from the source model prediction suggests the model has been adapted to a very specific target domain and further adapting towards one direction would increase the risk of failing to adapt to the continually changing distribution.
$\mathcal{R}_{anc}(x_{i};\theta^{*}_{m})=\frac{1}{N_{b}}\sum\limits_{x_{i},y_{i}\in{D}^{l}}||h(x_{i};\theta_{0})-h(x_{i};\theta^{*}_{m})||$
Smoothed Scoring for Model Selection: To determine the optimal model to select, we combine the cross-entropy loss and the anchor deviation as the final score. Direct summing the two metrics as the model selection score is sub-optimal due to the relative scale between the two metrics. Therefore, we propose to first normalize the two metrics. After the normalization, the relative size relationships of different candidate model scores are preserved, and a comparable validation loss can be conveniently obtained by addition, as below.
$\begin{split}&\forall m=1\cdots|\Omega|,\quad\;\mathrm{Norm}(x)=({x-\mathrm{min}(x)})/({\mathrm{max}(x)-\mathrm{min}(x)}),\quad S_{ce}(x_{i};\theta_{m}^{*})=\mathrm{Norm}(\mathcal{H}_{ce}(x_{i};\theta^{*}_{m}))\\ &S_{anc}(x_{i};\theta^{*}_{m})=\mathrm{Norm}(\mathcal{R}_{anc}(x_{i};\theta^{*}_{m})),\quad\mathcal{L}_{val}(x_{i};\theta^{*}_{m})=S_{ce}(x_{i};\theta^{*}_{m})+S_{anc}(x_{i};\theta^{*}_{m})\end{split}$
Considering that testing data stream is often highly correlated, i.e. temporally adjacent samples are likely subject to the same distribution shift, the optimal hyper-parameter/model should change smoothly as well. To encourage a smooth transition of the optimal model, we select the optimal model following an exponential moving average fashion as in Eq. 6 where we denote the moving average validation loss at the $t$-th batch as ¯$\bar{\mathcal{L}}_{val}^{t}$.
¯$\begin{split}&\bar{\mathcal{L}}_{val}^{t}(\theta_{m}^{*})=\beta\bar{\mathcal{L}}_{val}^{t-1}(\theta_{m}^{*})+(1-\beta)\frac{1}{|{B}_{t}^{l}|}\sum_{x_{j},y_{j}\in{B}_{t}^{l}}\mathcal{L}_{val}(h(x_{j};\theta_{m}^{*}),y_{j}),\quad\theta^{*}=\arg\min_{m}\bar{\mathcal{L}}_{val}^{t}(\theta_{m}^{*})\end{split}$
3.4 Sample Selection for HILTTA
The sample selection strategy should balance the objectives of active learning and model selection for more effective HILTTA. We achieve this target from two perspectives. First, the selected samples should prioritize low-confidence samples for more effective active learning, as adopted by SimATTA [Gui et al.(2024)Gui, Li, and Ji] through incremental K-means. However, selecting redundant low-confidence samples may harm the effectiveness of model validation. This motivates us to select samples approximating the testing data distribution, similar to the objective of active testing [Kossen et al.(2021)Kossen, Farquhar, Gal, and Rainforth]. Biasing towards either criterion would result in suboptimal HILTTA.
The above objectives could be achieved by a recent active learning approach, BADGE [Ash et al.(2020)Ash, Zhang, Krishnamurthy, Langford, and Agarwal], which takes the gradient of loss w.r.t. the penultimate layer feature as the uncertainty weighted feature as $g_{x}=\frac{\partial}{\partial\theta_{\theta_{out}}}\mathcal{L}_{ce}(h(x;\theta),\hat{y})$, on which K-Means++ is applied for active selection. The feature dimension is the product of the number of classes $C$ and penultimate layer channels $D$, i.e. $g_{x}\in\mathbbm{R}^{C\times D}$. Accumulating the feature over all categories results in overly high dimension features for large-scale classification tasks, e.g. $C=1000$ for ImageNet [Deng et al.(2009)Deng, Dong, Socher, Li, Li, and Fei-Fei]. To mitigate the ultra-high dimension challenge, we integrate uncertainty into feature representation by multiplying the uncertainty, measured by the margin confidence, to the feature representation as in Eq. 7, where $\mathrm{sort}^{d}$ refers to the descending sorting.
$\begin{split}&\hat{p}=\mathrm{sort}^{d}[h_{1}(x_{i};\theta),h_{2}(x_{i};\theta),\cdots h_{C}(x_{i};\theta)],&g_{i}=(1-\hat{p}_{1}+\hat{p}_{2})f(x_{i};\theta)\end{split}$
The above design prioritizes the difference between the probability of the most confident and second most confident classes as a quantification of model uncertainty.
Sample Selection via K-Margin: The samples to be selected for annotation are eventually obtained by conducting a K-Center clustering on the uncertainty-weighted features within each minibatch, as in Eq. 8, which is referred to as K-Margin clustering. We select K samples ${B}_{t}^{l}$ such that the maximal distance of all testing sample embedding is minimized. We deliberately exclude the selected samples in the previous minibatch, as opposed to SimATTA [Gui et al.(2024)Gui, Li, and Ji], because the labeled samples will serve as the validation set, and excluding new samples that may overlap with previously selected samples could undermine the effectiveness of model selection. The following problem can be solved via a greedy selection algorithm [Sener and Savarese(2018)].
$\min_{{B}_{t}^{l}\subseteq\{g_{i}\}_{i=1}^{N_{b}}}\max_{g_{i}\in\{g_{i}\}_{i=1}^{N_{b}}}\min_{c_{k}\in{B}_{t}^{l}}||g_{i}-c_{k}||,\quad s.t.\;|{B}_{t}^{l}|\leq K$ (8)
4 Experiments
4.1 Experimental setting
Datasets: We select four datasets for evaluation. The CIFAR10-C and CIFAR100-C [Hendrycks and Dietterich(2018)] are small-scale corruption datasets, with 15 different common corruptions, each containing 10,000 corrupt images with 10/100 categories. For our evaluation in large-scale datasets, we opt for ImageNet-C [Hendrycks and Dietterich(2018)], which also contains 15 different corruptions, each with 50,000 corrupt images in 1000 categories. Additionally, ImageNet-D [Rusak et al.(2022)Rusak, Schneider, Gehler, Bringmann, Brendel, and Bethge] is a style-transfer dataset, offering 6 domain shifts, each consisting of 10,000 images selected in 109 classes.
Implementation Details: We follow the continual TTA setting [Wang et al.(2022)Wang, Fink, Van Gool, and Dai, Niu et al.(2022)Niu, Wu, Zhang, Chen, Zheng, Zhao, and Tan] to evaluate TTA performance under long-term adaptation, where the target domain undergoes continuous changes. For CIFAR10-C and CIFAR100-C, we use a batch size of 200 and the annotation percentage of $3\%$, with respectively employed pre-trained WideResNet-28 [Zagoruyko and Komodakis(2016)] and ResNeXt-29 [Xie et al.(2017)Xie, Girshick, Dollár, Tu, and He] models. For ImageNet-C and ImageNet-D, we use a batch size of 64 for adaptation, and the annotation percentage is 3.2% of the total testing samples, with the ResNet-50 [He et al.(2016)He, Zhang, Ren, and Sun] pretrained model. For unsupervised training, we adopt the optimizer for the respective methods presented in the original paper. For supervised training, we use Adam optimizer, with a 1e-5 learning rate for all methods throughout all datasets. We consider seven distinct values for each hyper-parameter category regarding model selection. We defer more detailed experimental settings to the appendix.
4.2 Evaluations on HILTTA
Evaluation of Active Learning & Model Selection Strategies: We report the classification error on the four datasets in Tab. 1. Each method is specified by two attributes. The “Active Learning” attribute indicates the sample selection method for additional human annotation. The “Model Selection” attribute indicates whether the method enables hyper-parameter selection proposed in Sec. 3.2. We use TENT [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell] as an off-the-shelf TTA method. For methods that do not enable ”Model Selection”, we report three results, the worst (Worst), average (Avg.), and best (Best). The worst or best results are obtained by a fixed hyper-parameter trained with which the model exhibits the worst or best overall accuracy. The average results are obtained by averaging the overall accuracy across the models trained upon all candidate hyper-parameters. We make the following observations from the results in Tab. 1. i) Without additional annotation (TENT) the worst and average performance could vary substantially (best 62.96% v.s. worst 95.23% on ImageNet-C). The results suggest the sensitivity to hyper-parameter remains a major challenge. ii) With additional annotation (TENT w/ HIL), we first notice that applying active learning alone consistently improves over fully unsupervised adaptation in terms of worst, average and best results. iii) When model selection is further applied, the hyper-parameter is automatically chosen, resulting in a single error rate. We witness that for all active learning methods, our model selection strategy achieves significant improvement from the average error rate. In many cases, we could even surpass the best result. iv) Finally, our proposed K-Margin selection strategy combined with model selection almost outperforms all other active learning techniques, suggesting the effectiveness of balancing active learning and model selection for HILTTA.
Active Model ImageNet-C ImageNet-D CIFAR100-C CIFAR10-C
Learning Select. Worst Avg. Best Worst Avg. Best Worst Avg. Best Worst Avg. Best
Source $\times$ $\times$ 82.02 58.98 46.44 43.51
TENT [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell] $\times$ $\times$ 95.23 73.13 62.96 83.54 59.82 52.93 94.26 50.69 32.75 71.22 27.77 18.14
$\times$ 95.04 67.08 (+6.05) 59.60 79.15 55.01 (+4.81) 49.59 94.26 50.70 (-0.01) 30.87 64.61 24.91 (+2.86) 16.84
Random 59.42 (+13.71) 49.56 (+10.26) 31.50 (+19.19) 17.80 (+9.97)
$\times$ 95.23 70.47 (+2.66) 60.79 82.39 56.27 (+3.55) 49.85 94.37 51.33 (-0.64) 30.18 29.46 18.45 (+9.32) 15.77
Entropy [Shannon(1948)] 60.45 (+12.68) 49.48 (+10.34) 32.78 (+17.91) 15.82 (+11.95)
$\times$ 94.49 69.90 (+3.23) 60.73 80.15 56.40 (+3.42) 50.83 94.56 53.08 (-2.39) 32.08 60.62 25.92 (+1.85) 17.81
K-Center [Sener and Savarese(2018)] 60.56 (+12.57) 50.60 (+9.22) 33.25 (+17.44) 17.97 (+9.80)
TENT $\times$ 93.82 71.51 (+1.62) 62.78 79.06 54.91 (+4.91) 49.65 94.34 51.20 (-0.51) 31.59 59.06 24.41 (+3.36) 17.15
w/ HIL VeSSAL [Saran et al.(2023)Saran, Yousefi, Krishnamurthy, Langford, and Ash] 62.37 (+10.76) 49.56 (+10.26) 32.89 (+17.80) 17.81 (+9.96)
$\times$ 93.57 68.85 (+4.28) 60.66 78.32 54.25 (+5.57) 49.35 94.35 48.19 (+2.50) 30.53 70.23 25.36 (+2.41) 17.04
ASE [Kossen et al.(2022)Kossen, Farquhar, Gal, and Rainforth] 60.52 (+12.61) 53.08 (+6.74) 31.16 (+17.53) 17.49 (+10.28)
$\times$ 94.59 69.09 (+4.04) 60.71 79.99 55.25 (+4.57) 49.89 94.64 48.70 (+1.99) 30.36 31.28 19.35 (+8.42) 16.39
SimATTA [Gui et al.(2024)Gui, Li, and Ji] 60.46 (+12.67) 49.59 (+10.23) 31.23 (+19.46) 16.95 (+10.82)
K-Margin $\times$ 94.98 68.13 (+5.00) 58.49 76.27 53.77 (+6.05) 48.55 94.75 49.37 (+1.32) 30.11 27.81 18.21 (+9.56) 15.75
(OURS) 58.35 (+14.78) 48.74 (+11.08) 30.53 (+20.16) 15.87 (+11.90)
Table 1: Evaluation under HILTTA with average classification error (lower is better). ’Worst’, ’Avg.’, and ’Best’ denote the worst, average, and best performances, respectively, with the candidate hyper-parameter pool $\Omega$. The performance improvement by Active Learning and Model Selection is denoted within brackets as ($\delta$).
Evaluation under Alternative TTA Methods: Moreover, we investigate various off-the-shelf TTA methods in Tab. 2. Employing our proposed K-Margin as the sample selection method, we consistently observe enhancements for all off-the-shelf TTA methods when paired with our HIL model selection and learning strategies. These findings advocate for our active learning and model selection approach regardless of TTA methods. Importantly, even the simplest baseline, TENT, achieves comparable results to state-of-the-art TTA methods under HILTTA. The results suggest exploiting small amount of labeled data is very effective for TTA and the margin between different TTA methods are smaller than expected when equipped with human-in-the-loop.
TTA A.L. M.S. ImageNet-C ImageNet-D CIFAR100-C CIFAR10-C
Methods Worst Avg. Best Worst Avg. Best Worst Avg. Best Worst Avg. Best
$\times$ $\times$ 95.23 73.13 62.96 83.54 59.82 52.93 94.26 50.69 32.75 71.22 27.77 18.14
TENT [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell] $\times$ 94.98 68.13 (+5.00) 58.49 76.27 53.77 (+6.05) 48.55 94.75 49.37 (+1.32) 30.11 27.81 18.21 (+9.56) 15.75
58.35 (+14.78) 48.74 (+11.08) 30.53 (+20.16) 15.87 (+11.90)
$\times$ $\times$ 89.78 81.48 67.75 63.06 56.59 52.45 37.31 34.67 32.76 19.95 19.11 18.54
PL [Lee et al.(2013)] $\times$ 79.14 70.30 (+11.18) 62.15 50.93 49.25 (+7.34) 48.15 32.08 30.85 (+3.82) 29.93 17.02 16.67 (+2.44) 16.45
61.60 (+19.88) 49.01 (+7.58) 30.82 (+3.85) 16.97 (+2.14)
$\times$ $\times$ 99.18 79.56 64.68 91.51 64.51 53.69 90.72 43.89 32.90 52.87 23.22 17.26
SHOT [Liang et al.(2020)Liang, Hu, and Feng] $\times$ 99.29 74.49 (+5.07) 59.30 90.86 58.19 (+6.32) 48.37 87.30 41.05 (+2.84) 30.51 21.27 16.68 (+6.54) 15.66
60.66 (+18.90) 48.98 (+15.53) 31.13 (+12.76) 15.78 (+7.44)
$\times$ $\times$ 99.81 70.85 57.92 94.68 60.89 52.00 44.89 34.57 31.32 39.28 21.66 17.14
EATA [Niu et al.(2022)Niu, Wu, Zhang, Chen, Zheng, Zhao, and Tan] $\times$ 99.66 68.36 (+2.49) 54.37 94.75 56.93 (+3.96) 47.88 42.10 32.42 (+2.15) 29.51 25.06 17.82 (+3.84) 15.69
56.75 (+14.10) 49.17 (+11.72) 30.43 (+4.14) 15.79 (+5.87)
$\times$ $\times$ 90.60 70.70 62.20 56.19 53.37 51.72 39.00 33.84 31.90 20.22 19.10 17.45
SAR [Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan] $\times$ 89.03 64.81 (+5.89) 58.22 49.79 48.64 (+4.73) 47.50 32.24 30.63 (+3.21) 29.67 17.14 16.60 (+2.50) 15.63
58.91 (+11.79) 50.02 (+3.35) 30.00 (+3.84) 16.87 (+2.23)
$\times$ $\times$ 98.77 72.28 60.45 88.80 58.92 50.14 38.35 32.18 30.16 29.36 19.75 16.66
RMT [Döbler et al.(2023)Döbler, Marsden, and Yang] $\times$ 98.50 69.87 (+2.41) 56.67 87.96 55.51 (+3.41) 46.12 37.68 31.26 (+0.92) 29.02 27.04 18.35 (+1.40) 15.80
58.26 (+14.02) 46.28 (+12.64) 29.41 (+2.77) 16.45 (+3.30)
Table 2: Evaluation of different off-the-shelf TTA methods with K-Margin selection method under HILTTA protocol.
4.3 Ablation & Additional Study
Unveiling the Impact of Individual Components: We conduct a comprehensive ablation analysis using TENT as the off-the-shelf TTA method in Tab. 3. Compared with not adapting to testing data stream, fine-tuning with only the unsupervised training loss (Unsup. Train) could improve the performance, as expected. When human annotation is introduced, using the cross-entropy loss $\mathcal{H}_{ce}$ as validation objective (CE Valid.) does not guarantee a consistent improvement. As discussed in [Zhao et al.(2023)Zhao, Liu, Alahi, and Lin], this can be attributed to the challenges posed by the min-max equilibrium optimization problem across time, resulting in a notable performance decline. Incorporating additional anchor regularization for model selection (referred to as Anchor Reg.) consistently improves performance beyond the average error rate of the original TENT. This underscores the necessity of stronger regularization for model selection to achieve overall good TTA performance. Additionally, integrating EMA smoothing (EMA Smooth.) of validation loss and supervised training with labeled data (referred to as Super. Train) both contribute positively to enhancing TTA performance.
Remark Unsup. CE Anchor EMA Super. Error Rate (%) $\downarrow$
Train Valid. Reg. Smooth Train ImageNet-C ImageNet-D CIFAR100-C CIFAR10-C
Source - - - - - 82.02 58.98 46.44 43.51
TENT - - - - 73.13 59.82 54.97 27.77
w/ HIL - - - 90.52 62.66 67.82 21.36
w/ HIL - - 65.95 52.87 47.55 17.79
w/ HIL - 62.61 52.61 34.18 17.79
w/ HIL 58.35 48.74 30.53 15.87
Table 3: Ablation study of HILTTA with TENT  [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell] as off-the-shelf TTA method.
Strategy Error Rate (%) $\downarrow$
ImageNet-C ImageNet-D CIFAR100-C CIFAR10-C Avg.
Fixed Worst 94.98 76.27 94.75 27.81 73.45
Fixed Best* 58.49 48.55 30.11 15.75 38.23
Entropy [Morerio et al.(2018)Morerio, Cavazza, and Murino] 94.65 73.95 94.11 26.66 72.34
InfoMax [Musgrave et al.(2022)Musgrave, Belongie, and Lim] 94.56 77.00 93.14 28.81 73.38
MixVal [Hu et al.(2023)Hu, Liang, Liew, Xue, Bai, and Wang] 92.98 68.20 92.23 19.64 68.26
PitTTA [Zhao et al.(2023)Zhao, Liu, Alahi, and Lin] 60.75 47.95 47.99 16.23 43.23
HILTTA (Ours) 58.35 48.74 30.53 15.87 38.37
Table 4: Comparison of different model selection strategies. * indicate upper bound with fixed hyper-parameter
Evaluation of Alternative Model Selection Strategies: We further compare our approach with existing model selection strategies. Entropy [Morerio et al.(2018)Morerio, Cavazza, and Murino], InfoMax [Musgrave et al.(2022)Musgrave, Belongie, and Lim], and MixVal [Hu et al.(2023)Hu, Liang, Liew, Xue, Bai, and Wang] represent the state-of-the-art in model selection methods for Unsupervised Domain Adaptation. For the Oracle model selection strategy for online TTA, referred to as PitTTA [Zhao et al.(2023)Zhao, Liu, Alahi, and Lin], leverages all labels within the streaming test dataset $\mathcal{D}_{t}$ to conduct model selection based on accuracy comparison. The results in Tab. 4 reveal that our proposed methods consistently outperform alternative model selection approaches, yielding comparable or even superior performance when compared to the best outcomes achieved with fixed hyper-parameter choices.
Computation Efficiency: We conducted empirical studies to measure the wall-clock time of both the inference and adaptation steps in HILTTA with results in Tab. 5. In this setup, we only perform human-in-the-loop annotation for every N batches on the testing data stream. The inference time required is fixed for all methods and the adaptation time varies. We demonstrate that the additional model selection procedure significantly decreases the error rate (27.77% $\rightarrow$ 15.87%) with 9 times overall time required (1ms v.s. 9.1ms). More importantly, the overall time required can be further reduced by conducting sparser human annotation (larger N). With N=10, we witness a 10% absolute error rate decrease (27.77% $\rightarrow$ 17.13%) with only 2.8 times computation time. Considering annotating a single image would cost 3000 ms to 5000 ms, the adaptation time is considered to be negligible.
w/o HIL w/ HIL N=1 w/HIL N=3 w/HIL N=5 w/HIL N=10
Error Rate (%) 27.77 15.87 15.98 16.43 17.13
Inference Time (ms/sample) 0.40 0.40 0.40 0.40 0.40
Adaptation Time (ms/sample) 0.60 8.70 3.50 2.60 1.40
Table 5: Computation efficiency under different human intervention frequencies.
Figure 2: Performance of TENT[Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell] with and without HIL. Bold lines represent the performance with different hyper-parameter values, while dashed lines indicate the performance with model selection.
Detailed Analysis of Performance: We provide more detailed analysis against the selection of hyper-parameters. In particular, we investigate the performance of TENT with and without HIL and plot the results in Fig. 2. TENT w/o HIL and TENT w/o HIL (Average) refer to applying TENT without additional annotation and the average error rate respectively. TENT w/HIL Sup.Train (Random/Entropy/K-Margin) and TENT w/HIL Model Sel. (Random/Entropy/K-Margin) refer to augmenting TENT with supervised training on Random/Entropy/K-Margin labeled testing data and our proposed model selection respectively. The accuracy w/ HIL consistently surpasses all w/o HIL. Furthermore, random and entropy selection are consistently less effective than our proposed K-Margin selection for model selection purposes. Crucially, our final HILTTA approach consistently achieves significantly better performance than the worst hyper-parameter, and on ImageNet-C, we even surpass the model using the best fixed hyper-parameter.
5 Conclusion
In this study, we presented a novel perspective towards human-in-the-loop test-time adaptation (HILTTA) by synergizing active learning and model selection. Beyond active learning consideration, we further treat the labeled data as a validation set to enable the smooth selection of hyper-parameters and a subset for supervised training. we developed regularization techniques to augment the validation loss. A novel sample selection strategy is also proposed for HILTTA. Then we integrate HILTTA to multiple sample selection methods and off-the-shelf TTA methods and witness consistent improvement jointly brought by active learning and model selection over choosing the worst hyper-parameter and average performance. These explorations could help the community rethink the way to exploit the limited annotation budget in TTA tasks.
Appendix
In this appendix, we first provide a detailed description of our HILTTA algorithm. We further present additional insights and analysis including performance evolution, selecting two hyper-parameters, additional labeling budgets, computation complexities, and comparison with additional methods. Finally, we provide more details of experiments.
Appendix A Overall Algorithm for HILTTA
Input : Source Model $\theta_{0}$; Candidate Hyper-parameters $\Omega=\{\omega_{m}\}$; Testing Data Batches $\{{B}_{t}\}$
Output : Predictions $\hat{\mathcal{Y}}=\{\hat{y}_{i}\}$
for $t=1$ to $T$ do
       # Make Predictions:
       $\forall x_{i}\in{B}_{t},\;\hat{y}_{i}=h(x_{i};\theta^{t-1*}),\;\hat{\mathcal{Y}}=\hat{\mathcal{Y}}\cup\hat{y}_{i}$
       # Oracle Annotation:
       Select labeled subset ${B}_{t}^{l}$ by Eq. 8
       for $m=1$ to $N_{m}$ do
             # Unsupervised Model Adaptation:
             $\theta_{m}^{t*}=\arg\min\limits_{\theta}\frac{1}{N_{b}}\sum\limits_{x_{i}\in{B}_{t}^{u}}\mathcal{L}_{tr}^{u}(h(x_{i};\theta);\omega_{m})$
      # Update Moving Average Validation Loss:
       Update ¯$\bar{\mathcal{L}}_{val}^{t}$ by Eq. 6.
       # Model Selection by Eq. 6:
       $\theta^{t*}=\arg\min\limits_{m}\bar{\mathcal{L}}_{val}^{t}(\theta^{*}_{m})$
       # Supervised Model Adaptation:
       $\theta^{t*}=\arg\min_{\theta}\frac{1}{N_{b}}\sum\limits_{x_{i},y_{i}\in{B}_{t}^{l}}\mathcal{L}_{tr}^{l}(h(x_{i};\theta),y_{i};\omega_{m})$
return Predictions $\hat{\mathcal{Y}}$;
Algorithm 1 Human-in-the-Loop TTA
We introduce two types of training losses for model updates and a validation loss for model selection. In the model selection stage, multiple candidate models are obtained by training with an unsupervised TTA training loss, denoted as $\mathcal{L}_{tr}^{u}$. The exact design of $\mathcal{L}_{tr}^{u}$ follows the off-the-shelf TTA methods. Then we select the best candidate model $\theta^{t*}$ by $\arg\min\limits_{m}\bar{\mathcal{L}}_{val}^{t}(\theta^{*}_{m})$. After the model selection stage, we further update the model in a supervised manner by optimizing a cross-entropy loss, denoted as $\mathcal{L}_{tr}^{l}$. We finally present the overall algorithm for Human-in-the-Loop Test-Time Adaptation in Alg. 1.
Appendix B Further Insights and Analysis Findings
B.1 Temporal Evolution Analysis of HILTTA
We present an analysis of the evolution of selected hyper-parameters over time. As shown in Fig. 3, the hyper-parameters selected by our HILTTA model selection method always hover around the upper bound “Fixed Best” while competing methods deviate further away from the upper bound. The advantage is also reflected in the accumulated error plot. In particular, methods such as MixVal [Hu et al.(2023)Hu, Liang, Liew, Xue, Bai, and Wang], are susceptible to catastrophic failure as they struggle to adapt to continuously changing domains.
Figure 3: (Left) Average selected learning rate per corruption for TENT on ImageNet-C. (Right) Average error rate per corruption for TENT on ImageNet-C.
B.2 Selecting Two Hyperparameters
In certain cases, there are more than one hyper-parameter should be selected. We evaluate HILTTA combined with PL [Lee et al.(2013)] on CIFAR10-C by selecting two hyper-parameters, i.e. the self-training threshold and learning rate. As shown in Tab. 6, our HILTTA achieves the 3rd best result among 25 candidate hyper-parameter sets. The results suggest our method is agnostic to the number of hyper-parameters and can be utilized for real-world TTA tasks where an annotation budget is available.
LR/Threshold 0.1 0.2 0.4 0.6 0.8 AVG. Error HILTTA (OURS)
1e-5 17.17 17.12 17.19 17.22 17.22 28.96 16.41
1e-4 17.08 17.10 17.01 17.13 17.04
1e-3 16.99 16.76 16.54 16.49 16.43
1e-2 15.89 15.56 20.29 23.69 23.56
1e-1 22.41 81.89 88.64 88.89 88.78
Table 6: Select two hyperparameters at the same time with PL on Cifar10-C, error rate is reported.
B.3 Performance Under Different Labeling Budgets
We investigate the effectiveness of HILTTA under different annotation budgets. Specifically, we vary the annotation ratio from $0\%$ to $10\%$. As shown in Tab. 7, with the increase of annotation budget, we observe consistent improvement in error rate for all off-the-shelf TTA methods. Importantly, naive methods, e.g. TENT, exhibit superior performance at higher labeling budget than more sophisticated ones, e.g. EATA, SAR and RMT. This suggests that limited labeled data could substantially ease the challenges of TTA.
Budget TENT PL SHOT EATA SAR RMT
0% 27.77 19.11 23.22 19.57 19.10 19.75
1% 17.03 18.09 16.61 17.27 18.06 17.36
2% 16.35 17.30 16.21 16.81 17.30 16.41
3% 15.87 16.97 15.78 16.40 16.87 16.45
5% 15.13 16.45 15.44 15.83 16.12 16.78
10% 14.68 15.66 14.78 15.01 15.41 16.85
Table 7: Performance under different labeling budgets on CIFAR10-C, average error rates are reported.
B.4 Comparing Adaptation Time & Annotation Time
We carried out empirical studies by measuring the wall-clock time lapses of the inference step (Inference Time), adaptation step (Adaptation Time), and human annotation (Annotation Time) per sample. In specific, the "Inference Time" refers to the time-lapse of pure forward pass, the "Adaption Time" refers to the time-lapse of backpropagation and weights update, and "Annotation Time" refers to the estimated time lapse required for human to annotate images at a 3% labeling rate. We evaluate under three categories of annotation cost, i.e. 2 sec/sample, 5 sec/sample, and 10 sec/sample. As seen from Tab. 8, despite the adaptation time being higher than the inference time (8.7 ms v.s. 0.4 ms), the annotation time is the most expensive step. These results suggest model selection and adaptation do not slow down the throughput under the HILTTA protocol.
Inference Adaptation Annotation (2 sec/sample) Annotation (5 sec/sample) Annotation (10 sec/sample)
Time (ms/sample) 0.40 8.70 60.00 150.00 300.00
Table 8: Comparing HILTTA and annotation time consuming with a 3% label rate on CIFAR10-C
B.5 Computation Complexity of Competing Methods
We conclude that existing stream-based active learning (VeSSAL [Kossen et al.(2022)Kossen, Farquhar, Gal, and Rainforth]) and active TTA ( SimATTA [Gui et al.(2024)Gui, Li, and Ji]) are inherently more computationally expensive than HILTTA due to the following reasons. First, VeSSAL adopts gradient embedding, i.e. $g(x_{i})=\partial l(f(x_{i};\theta),y_{i})/\partial\theta_{L}$, where $\theta_{L}\in\mathbb{R}^{K\times D}$ refers to the classifier weights. This results in a very high dimensional gradient embedding $g(x_{i})\in\mathbb{R}^{D\cdot K}$. The probability of labeling a sample $p_{i}$ involves calculating the inverse of covariance matrix, i.e. $p_{i}\propto g(x_{i})^{\top}\hat{\Sigma}^{-1}g(x_{i}),\;s.t.\hat{\Sigma}=\sum_{x_{i}\in\mathcal{B}}g(x_{i})g(x_{i})^{\top}\in\mathbb{R}^{DK\times DK}$. Hence, VeSSAL can hardly deal with classification tasks with a large number of classes, e.g. ImageNet, where $K$ is large and calculating a large matrix inverse is inherently time-consuming. Second, SimATTA employs a K-Means clustering algorithm for the selection of samples to annotate. Since K-Means is an iterative approach the computation cost of SimATTA is relatively high as well. In contrast, our proposed sample selection is built upon K-Center clustering which sequentially selects the sample with the highest distance to cluster centers (greedy algorithm), the computation cost is significantly lower than VeSSAL and SimATTA. An empirical study in Tab. 9 reveals that our method is at least 3 times faster than VeSSAL and SimATTA in Adaptation Time.
Methods Error Adaptation Time (ms/sample)
VeSSAL 24.41 26.90
SimATTA 17.66 22.50
HILTTA (OURS) 15.87 8.70
Table 9: Error rate and adaptation wall time comparing to different methods on CIFAR10-C dataset
B.6 Studies for Regularization in HILTTA
We argue that proposed anchor regularization and EMA smoothing for model selection in HILTTA prevent the model from overly adapting to local data distribution on the testing data stream. Removing them may help adaptation to a single domain but harm the performance of future domains. As shown in Tab. 10, in the continual TTA setting, model selection without regularization is much worse than the one with regularization (17.25% v.s. 15.87%), suggesting regularization is necessary for continual TTA.
Regularization Gaussian Shot Impulse Defocus Jpeg AVG.
Without 25.04 20.23 28.57 12.22 19.37 17.25
With 25.93 20.13 27.49 11.59 18.26 15.87 (+1.38)
Table 10: Comparison between w/ and w/o regularization for HILTTA with TENT on Cifar10-C. error rates are reported.
B.7 Comparing with Knowledge Distillation and Ensemble Learning
We carried out additional experiments by comparing with ensemble learning and multi-teacher knowledge distillation methods. Specifically, for ensemble learning, we average the posterior predicted by each candidate model and make a prediction with the average posterior. For multi-teacher knowledge distillation, we use the pseudo label predicted by the ensemble model for self-training. As observed from Tab. 11, both ensemble learning and multi-teacher knowledge distillation yield inferior results than HILTTA. This is most likely caused by incorporating poor candidate models (ensemble learning) and using poor pseudo labels for self-training, i.e. learning from noisy labeled data (multi-teacher knowledge distillation).
Method Average Performance on all candidate model Ensemble Learning Muli-teacher Knowledge DIstillation HILTTA (OURS)
Error 27.77 19.48 26.42 15.87
Table 11: Comparison with Ensemble Learning and Knowledge Distillation with TENT on CIFAR10-C.
B.8 Random Batchsizes
We recognize that in certain scenarios, test data may arrive inconsistently and intermittently. Results from Tab. 12 demonstrate that our proposed HILTTA framework is adaptable to such varying batchsize. At each iteration, we vary the batchsize randomly among the choices [50, 100, 150, 200]. This versatility leads to even greater improvements compared to fixed batch size. This enhanced performance stems from the fact that the optimal hyperparameter values vary with different batch sizes, thus benefiting significantly from our adaptive model selection approach.
Batchsize w/o HIL HILTTA
200 27.77 15.87 (+11.90)
Random choose from [50,100,150,200] 35.89 16.39 (+19.50)
Table 12: Adaptation error with various batchsize with TENT on CIFAR10-C.
B.9 Ablation Study with RMT
We have conducted wider experiments to analyze the ablation study with RMT [Döbler et al.(2023)Döbler, Marsden, and Yang] as the off-the-shelf TTA method in Tab. 13 exhibits a consistent trend with the findings presented in the manuscript, providing additional support for our analysis and demonstrating the effectiveness of our proposed approach.
Remark Unsup. CE Anchor EMA Super. Error (%) $\downarrow$
Train. Valid. regular. Smooth. Train. ImageNet-C ImageNet-D CIFAR100-C CIFAR10-C
Source - - - - - 82.02 58.98 46.44 43.51
RMT - - - - 72.28 58.92 32.18 19.75
w/ HIL - - - 98.34 49.98 36.28 24.94
w/ HIL - - 67.98 51.40 31.53 19.18
w/ HIL - 61.33 50.23 30.75 17.33
w/ HIL 58.26 46.28 29.41 16.45
Table 13: Ablation study of HILTTA. Average classification error is reported for combing with RMT.
Appendix C Experiment Details
C.1 Detailed Settings for Datasets
Following prior works [Döbler et al.(2023)Döbler, Marsden, and Yang, Wang et al.(2022)Wang, Fink, Van Gool, and Dai, Niu et al.(2022)Niu, Wu, Zhang, Chen, Zheng, Zhao, and Tan], we perform continual test time adaptation in the following datasets, detailed in follows.
CIFAR10-C [Hendrycks and Dietterich(2018)] is a small-scale corruption dataset, with 15 different common corruptions, each containing 10,000 corrupt images of dimension (3, 32, 32) with 10 categories. We evaluate severity level 5 images, with "Gaussian $\rightarrow$ shot $\rightarrow$ impulse $\rightarrow$ defocus $\rightarrow$ glass $\rightarrow$ motion $\rightarrow$ zoom $\rightarrow$ snow $\rightarrow$ frost $\rightarrow$ fog $\rightarrow$ brightness $\rightarrow$ contrast $\rightarrow$ elastic $\rightarrow$ pixelate $\rightarrow$ jpeg" sequence.
CIFAR100-C [Hendrycks and Dietterich(2018)] is a small-scale corruption dataset, with 15 different common corruptions, each containing 10,000 corrupt images of dimension (3, 32, 32) with 100 categories. We evaluate severity level 5 images, with "Gaussian $\rightarrow$ shot $\rightarrow$ impulse $\rightarrow$ defocus $\rightarrow$ glass $\rightarrow$ motion $\rightarrow$ zoom $\rightarrow$ snow $\rightarrow$ frost $\rightarrow$ fog $\rightarrow$ brightness $\rightarrow$ contrast $\rightarrow$ elastic $\rightarrow$ pixelate $\rightarrow$ jpeg" sequence.
ImageNet-C [Hendrycks and Dietterich(2018)] is a large-scale corruption dataset, with 15 different common corruptions, each containing 50,000 corrupt images of dimension (3, 224, 224) with 1000 categories. We evaluate the first 5,000 samples with severity level 5 following [Croce et al.(2021)Croce, Andriushchenko, Sehwag, Debenedetti, Flammarion, Chiang, Mittal, and Hein]. The domain sequence is "Gaussian $\rightarrow$ shot $\rightarrow$ impulse $\rightarrow$ defocus $\rightarrow$ glass $\rightarrow$ motion $\rightarrow$ zoom $\rightarrow$ snow $\rightarrow$ frost $\rightarrow$ fog $\rightarrow$ brightness $\rightarrow$ contrast $\rightarrow$ elastic $\rightarrow$ pixelate $\rightarrow$ jpeg".
ImageNet-D [Rusak et al.(2022)Rusak, Schneider, Gehler, Bringmann, Brendel, and Bethge], as a large-scale style-transfer dataset, is built upon DomainNet [Peng et al.(2019)Peng, Bai, Xia, Huang, Saenko, and Wang], which encompasses 6 domain shifts (clipart, infograph, painting, quickdraw, real, and sketch). The dataset focuses on samples belonging to the 164 classes that overlap with ImageNet. Following [Marsden et al.(2024)Marsden, Döbler, and Yang], we specifically select all classes with a one-to-one mapping from DomainNet to ImageNet, resulting in a total of 109 classes and excluding the quickdraw domain due to the challenge of attributing many examples to a specific class. After processing, it contains 5 domain shifts with each 10,000 images selected in 109 classes. The domain sequence is "clipart $\rightarrow$ infograph $\rightarrow$ painting $\rightarrow$ real $\rightarrow$ sketch".
C.2 Detailed Settings for Comparing Methods
We mainly follow the official implementation of each method and use the same optimization strategy. But in certain cases, such as transferring from another protocol, we make minor changes to the configuration.
Source serves as the baseline for inference performance without any adaptation.
TENT [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell]. It focuses on updating batch normalization layers through entropy minimization. We follow the official implementation [^1] https://github.com/DequanWang/tent of TENT to update BN parameters.
PL [Lee et al.(2013)]. We implement self-supervised training on unlabeled data updating all parameters by cross-entropy with an entropy threshold $E=\theta\times log(\#class)$. For CIFAR10-C and CIFAR100-C, we use an SGD optimizer with a 1e-3 learning rate, while an SGD optimizer with a 2.5e-5 learning rate for the rest.
SHOT [Liang et al.(2020)Liang, Hu, and Feng]. It freezes the linear classifier and trains the feature extractor by balancing prediction category distribution, coupled with pseudo-label-based self-training. We follow its official code [^2] https://github.com/tim-learn/SHOT to update the feature extractor but remove the threshold. For CIFAR10-C and CIFAR100-C, we use an SGD optimizer with a 1e-3 learning rate, while an SGD optimizer with a 2.5e-5 learning rate.
EATA [Niu et al.(2022)Niu, Wu, Zhang, Chen, Zheng, Zhao, and Tan]. It adapts batch normalization layers using entropy minimization but introduces an additional fisher regularization term to prevent drastic parameter changes. We follow the official implementation [^3] https://github.com/mr-eggplant/EATA of EATA to update BN parameters. We hold 2000 fisher training samples and keep the entropy threshold $E_{0}=0.4\times log(\#class)$.
SAR [Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan]. It updates batch normalization layers with a sharpness-aware optimization approach. We mainly follow its official code [^4] https://github.com/mr-eggplant/SAR, but set $e_{0}=0$ for CIFAR10-C and CIFAR100-C, preventing failure cases of frequently resetting the model’s weight.
RMT [Döbler et al.(2023)Döbler, Marsden, and Yang]. It adapts in a teacher-student mode, incorporating cross-entropy consistency loss and contrastive loss. We follow the official implementation [^5] https://github.com/mariodoebler/test-time-adaptation.
ASE [Kossen et al.(2022)Kossen, Farquhar, Gal, and Rainforth]. It uses surrogate estimators to sample in a uncertainty distribution to sample with low bias. We follow its official implementation [^6] https://github.com/jlko/active-surrogate-estimators.
SIMATTA [Gui et al.(2024)Gui, Li, and Ji]. As an active test-time adaptation method, it selects labeled samples through incremental clustering and updates the model by minimizing cross-entropy. To reduce the computational and memory cost of incremental clustering, we keep the maximum length of anchors to 50 in all experiments. For CIFAR10-C and CIFAR100-C, we set the lower entropy bound $e_{l}$ to 0.005 and 0.01 for the higher entropy bound $e_{u}$. While for ImageNet-C and ImageNet-D, we set the lower entropy bound $e_{l}$ to 0.2 and 0.4 for the higher entropy bound $e_{u}$.
VeSSAL [Saran et al.(2023)Saran, Yousefi, Krishnamurthy, Langford, and Ash]. As a stream-based active learning method, it uses volume sampling for streaming active learning. We also follow its official implementation [^7] https://github.com/asaran/VeSSAL, using feature embedding for stream-based data selection.
C.3 Hyper-parameter Choices for Model Selection
To achieve a fair and diverse validation process, we select three main kinds of hyper-parameters, i.e. threshold, loss coefficient, and learning rate. Detailed hyper-parameter settings along with the search space are available in Tab. 14.
TTA method Hyper-parameter Search Space
TENT [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell] Learning Rate $\eta$5e-5, 1e-4, 2.5e-4, 5e-4, 1e-3, 2.5e-3, 5e-3}
PL [Lee et al.(2013)] Entropy Threshold $\tau$0.05, 0.1, 0.2, 0.4, 0.6, 0.8, 1.0}
SHOT [Liang et al.(2020)Liang, Hu, and Feng] Loss Coefficient $\beta$0.1, 0.25, 0.5, 1.0, 2.5, 5.0, 10.0}
EATA [Niu et al.(2022)Niu, Wu, Zhang, Chen, Zheng, Zhao, and Tan] Learning Rate $\eta$5e-5, 1e-4, 2.5e-4, 5e-4, 1e-3, 2.5e-3, 5e-3}
SAR [Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan] Learning Rate $\eta$5e-5, 1e-4, 2.5e-4, 5e-4, 1e-3, 2.5e-3, 5e-3}
RMT [Döbler et al.(2023)Döbler, Marsden, and Yang] Loss Coefficient $\lambda_{CL}$0.1, 0.25, 0.5, 1.0, 2.5, 5.0, 10.0}
Table 14: Overview of the TTA methods validated and their associated hyper-parameters